OpenStudy (anonymous):

For the following system, use the second equation to make a substitution for y in the first equation.

OpenStudy (anonymous):

3x + y = 2 y + 4 = 3x

OpenStudy (jack1):

EQN 1: 3x + y = 2 EQN 2: y + 4 = 3x use the second equation to make a substitution for y in the first equation. so for ENQ 2: rearrange it so it is in the form y = ...something then put the value for ...something into eqn 1 where there is a y ie if your 2 equations were: EQN 1: 6x + 2y = 28 EQN2: y + 25 = 10x you would rearrange equation 2 so it was in the form y = ...something so y + 25 = 10x => y = 10x - 25 then put the value for ...something into eqn 1 where there is a y so 6x + 2y = 28 => 6x + 2(10x - 25) = 28 6x + 20x - 50 = 28 26x -50 = 28 26x = 78 x = 78/26 x = 3

OpenStudy (jack1):

does that make sense @chronicles ?

OpenStudy (anonymous):

yes thanks

OpenStudy (jack1):

sweet all good